OpenStudy (anonymous):

please help! I did the questions but I need someone help me check them? Fan, medal, testimony

OpenStudy (anonymous):

I got C

OpenStudy (anonymous):

we can do this, you need to find two distances, from one wall to the source and from the other wall to the source you supposed to use trig?

OpenStudy (anonymous):

iirc, the ratio for a 30-60-90 is 1:rt(3):2

OpenStudy (anonymous):

@satellite73 I think so yeah. this in my geometry class but in this current assignment I have trig problems

OpenStudy (anonymous):



OpenStudy (anonymous):

C is correct. By solving , I got 180/sqrt(3) which is 60sqrt(3)= 103.92

OpenStudy (anonymous):

\[\tan(60)=\frac{60}{x}\iff x=\frac{60}{\tan(60)}\]\[\tan(30)=\frac{40}{y}\iff y=\frac{40}{\tan(30)}\]

OpenStudy (anonymous):

you don't really need trig for these you can use the triangle ratios too as done above

OpenStudy (anonymous):

oh okay, so I was right?

OpenStudy (anonymous):

what is your answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

actually you don't have to do any work to solve this 95% of the time the answer is C, and it is C here too

OpenStudy (anonymous):

I got D